OpenStudy (anonymous):

(a/(a^(2)+a-2))+(2/(a^(2)-5a+4))

OpenStudy (anonymous):

Simplify this expression

OpenStudy (campbell_st):

if you factorise the denominators then you have \[\frac{a}{(a +2)(a - 1)} + \frac{2}{(a - 4)(a - 1)}\] so the common denominator will be (a +2)(a -1)(a -4) so the numerator in the 1st fraction needs to me multiplied by (a -4) the numerator in the 2nd fraction needs to be multiplied by (a + 2) so \[\frac{a(a -4)}{(a+2)(a -1)(a -4)} + \frac{2(a + 2)}{(a+2)(a -1)(a -4)}\] so if you colelct like terms in the numerator and then factorise it I think there will be a common factor that can be cancelled.